# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ 5 IGB: emergence and amplification
By leveraging Theorem 4.1, we are able to theoretically estimate the p.d.f. of $f_{0}$, building upon the distributions defined in Eq. (5) and Eq. (6). These estimations, described in App. E.1 and App. G.3.1, enable us to systematically evaluate $p_{f_{0}}\left({x}\right)$ across a spectrum of model designs.
@@
    $p_{\Delta_{O}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\Delta_{\mu},2\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right)\,.$ (11)
@@
    Substituting Eq. (11) into Eq. (4) allows us to invert the equation linking $f_{0}$ to $\Delta_{\mu}$, thereby expressing $\Delta_{\mu}$ as a function of $f_{0}$.
@@
    $p_{\Delta_{\mu}}\left({y}\right)=\mathcal{N}\left(y;0,2\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\right).$ (12)
@@
    Starting from Eq. (12) and the relationship $\Delta_{\mu}(f_{0})$ from point (B), we obtain $p_{f_{0}}\left({x}\right)$ through a change of variables, that is, by applying the formula:
    $p_{f_{0}}\left({x}\right)\,dx=p_{\Delta_{\mu}}\left({y(x)}\right)\,dy\,.$ (13)
This analysis not only identifies the architectural elements critical to the emergence of IGB but also quantifies their influence. [^1] We note that, to derive $p_{f_{0}}\left({x}\right)$, we work in the infinite-width limit. This is however not a necessary condition for the appearance of IGB (see e.g. experiments with finite-width models in App. I), but rather an assumption that allows us to derive the theoretical curves.
@@ 5.1 Single hidden layer
Figure 5: The distribution $p_{f_{0}}\left({x}\right)$ in a single-hidden-layer perceptron, for different choices of activation functions and with/without max pooling.
Let us begin by considering the case with $L=1$; in this scenario, the element of non-linearity in the network (such as the activation function, pooling layer, etc.) can give rise to IGB. In our analysis we consider three different setups to highlight the different behaviors induced by non-linearity elements. Fig. 5 shows $p_{f_{0}}\left({x}\right)$ (both empirical histogram and theoretical curves) for the following cases:
  • Linear: $p_{f_{0}}\left({x}\right)$ asymptotically converges to a delta distribution peaked on $f_{0}=1/2$. [^2] By asymptotically, we mean in the $D\rightarrow\infty$} limit. The theoretical curve shown in the plot takes into account the finite dataset size effects [^3] To take finite dataset-size effects into account, we substitute the population data distribution, with the empirical one, defined over the finite set of $D$ points. (since in real simulations $D<\infty$) as discussed in App. D.
  • ReLU: In this case, $p_{f_{0}}\left({x}\right)$ does not concentrate at $f_{0}=1/2$ and stays asymptotically wide (we detail this in App. D, Fig. D.1), so $f_{0}$ will, w.h.p., be away from 1/2. The mode of the distribution remains, as for the linear case, at $f_{0}=1/2$. Yet, in this case, the fluctuations from the peak are not due to finite size effects and remain finite in the limit of infinite data points.
@@
    $\rho^{(1;m)}_{l}$ $=\rho^{(m)}\left({\left\{g\left(h^{(1)}_{j}\right)\right\}_{j\in S_{l}^{m}}}\right)\,$, (14)
    where, in our notation, $\rho^{(m)}\left({\cdot}\right)$ indicates a pooling function with kernel size equals to $m$, $S_{l}^{m}$ indicate the $l^{\mathrm{th}}$ subgroup of $m$ nodes. In particular, for the max pooling
    $\rho^{(1;m)}_{l}\equiv\max_{j\in S_{l}^{m}}\left\{\max(0,g^{(1)}_{j})\right\}$. (15)
@@
Note how, for the output layer, permutation symmetry in nodes corresponds to symmetry between classes. The latter is always preserved at the ensemble level, [^4] Here, ”ensemble” refers to all possible random initializations specified by the chosen scheme (e.g., Kaiming Normal). in fact in all three cases $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$, where $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu\equiv\int p_{\mathcal{W}}\left({\tilde{\mathcal{W}}}\right)f_{0}\left(\tilde{\mathcal{W}}\right)d\tilde{\mathcal{W}}$ indicate the average of $f_{0}\left(\mathcal{W}\right)$ over the ensemble of weight initializations. However, while in the linear case, the symmetry between classes is also conserved on the single element of the ensemble, [^5] In this case, we have $\lim_{D\rightarrow\infty}f_{0}\left(\mathcal{W}\right)=\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$. in the other two cases the single realization diverges, w.h.p. from the mean over the ensemble. NSB thus results in a breaking of the symmetry between classes and a consequent self-averaging breakdown for the observable $f_{0}$. This difference between the two estimates is not due to finite network size effects; our predictions (theoretical curves in Fig. 5) are, in fact, formally exact in the limit of infinite size.
@@
Our results can be extended to a generic activation function. In particular, we can clarify (App. F) what is the fundamental attribute of the activated nodes, i.e. passed through the activation function and the pooling layer, $\{\rho^{(1;m)}_{i}\}_{i}$, that triggers NSB and consequently IGB:
@@
Consider a Gaussian distributed dataset processed through a single hidden layer perceptron, in the limit of infinite width hidden layer, and whose weights are initialized according to the Kaiming normal scheme (with null bias weights). Then, we will have an absence of IGB in the network if, and only if, for the generic $i^{\text{th}}$ node of the hidden layer: $\left\langle{\rho^{(1;m)}_{i}}\right\rangle=0\;,\;\;\forall i.$ (16) Or conversely, IGB will emerge if, and only if $\left\langle{\rho^{(1;m)}_{i}}\right\rangle\neq 0\;,\;\;\forall i.$ (17)
@@ 5.2 Conditions for strong IGB
Consider a centered Gaussian distributed dataset that undergoes preprocessing through the standardization $\psi\left(\boldsymbol{\xi^{(a)}}\right)=\boldsymbol{\xi^{(a)}}+\boldsymbol{K},$ where $\boldsymbol{K}$ is an offset with the same dimensionality of the input. The data is then processed through a Multi-Layer Perceptron (MLP) with $L$ hidden layers. This MLP employs ReLU nonlinearities and max pooling layers with a kernel size of $m$, and the weights are initialized using the Kaiming normal scheme (with zero biases). In the limit of infinite width for hidden layers, the following scenarios are observed: I. Shifting the center of the input distribution by a vector with the norm $|\boldsymbol{K}|$, even when $L=1$ and $m=1$, results in: $\lim_{|\boldsymbol{K}|\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (18) II. With an increase in the kernel size of the pooling layer, even for a single hidden layer ($L=1$), and in the absence of standardization offset ($|\boldsymbol{K}|=0$), one has: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (19) III. In the infinite-depth limit, even in the absence of pooling layers ($m=1$) and standardization ($|\boldsymbol{K}|=0$): $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (20)
@@ 5.3.3 Effects of IGB on Pre-Trained Models
Figure 6: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained EfficientNetV2 with the binary dataset (C10). The difference lies in the untrained head: a fully connected layer (PT-ENetV2-FC) versus a MLP (PT-ENetV2-MLP). More details in App. I.4. Increasing head depth amplifies IGB, consistently with our analysis.
@@ Appendix A Notation
$h^{(l+1)}_{i}$ $=\sum_{j}w^{(l+1)}_{ij}\rho^{(l;m)}_{j}\,$, (21)
$g^{(l)}_{i}$ $=g\left({h^{(l)}_{i}}\right)\,$, (22)
$\rho^{(l;m)}_{i}$ $=\rho^{(m)}\left({\left\{g\left(h^{(l)}_{j}\right)\right\}_{j\in S_{l}^{m}}}\right),$ (23)
where $l\in\{0,\dots,L+1\}$, $g\left({\cdot}\right),:\mathbb{R}\rightarrow\mathbb{R}$ represents the activation function, $\rho^{(m)}\left({\cdot}\right),:\mathbb{R}^{m}\rightarrow\mathbb{R}$ denotes the pooling function with kernel size equals to $m$, $S_{l}^{m}$ indicate the $l^{\mathrm{th}}$ subgroup of $m$ nodes; $\rho^{(0;m)}_{i}\equiv\boldsymbol{\xi}$ and $h^{(L+1)}_{c}\equiv O^{(c)}$. The principal notation is summarized in Fig. A.1.
@@
  • $F_{X}\left({x}\right)$:Given a r.v. $X$, we denote its cumulative distribution function (c.d.f.) as $F_{X}\left({x}\right)$, i.e., $F_{X}\left({x}\right)=\mathbb{P}\left(X<x\right)$. Considering the r.v. $X(\chi,\mathcal{W})$, which is a function of two independent sets of random variables $\chi$ and $\mathcal{W}$, when one of these sources of randomness is fixed, we explicitly indicate the active source in the notation. For example, $F_{X}^{(\chi)}\left({x}\right)=\mathbb{P}\left(X<x\mid\mathcal{W}\right)$.
  • $f_{c}$: fraction of dataset elements classified as belonging to class $c$. The argument $M$ indicates the total number of output nodes for the variable definition, i.e. the number of classes considered. For binary problems we omit this argument ($f_{0}$) as there is only one non-trivial possibility, i.e. $M=2$ .
  • $f_{\tilde{c}}$: the set $\left\{f_{c}\right\}_{c=0}^{M-1}$ contain the same elements of $\left\{f_{\tilde{c}}\right\}_{c=0}^{M-1}$, but these are ranked by magnitude, such that $f_{\tilde{0}}$ is the greatest output value between the $M$ elements of the set, $f_{\tilde{1}}$ the second one and so on.
  • $\boldsymbol{g}^{(l)}$: vector of the $l^{\text{th}}$ hidden layer nodes, $\left(g^{(l)}_{0},\dots,g^{(l)}_{N_{l}-1}\right)$, after passing through the activation function; $g^{(l)}_{i}$ indicate the component corresponding to node $i$.
  • $\boldsymbol{h}^{(l)}$: vector of the $l^{\text{th}}$ hidden layer nodes, $\left(h^{(l)}_{0},\dots,h^{(l)}_{N_{l}-1}\right)$, before passing through the activation function; $h^{(l)}_{i}$ indicate the component corresponding to node $i$.
@@
  • $\mathcal{N}\left(x;\mu,\sigma^{2}\right)$: Given a Gaussian r.v. $X$, we indicate with $\mathcal{N}\left(x;\mu,\sigma^{2}\right)$ the p.d.f. computed at $X=x$, i.e. $\mathcal{N}\left(x;\mu,\sigma^{2}\right)\equiv p_{X}\left({x}\right)=\frac{e^{-\frac{1}{2\sigma^{2}}(x-\mu)^{2}}}{\sqrt{2\pi\sigma^{2}}}$
  • $O^{(c)}_{M}$: output layer node; $c$ is the node index; the index M, instead, indicate the total number of nodes considered. For binary problems we omit the subscript index to keep the notation lighter, i.e. $O^{(c)}$ .
  • $O^{(\tilde{0})}_{M}$: the set $\left\{O^{(\tilde{c})}_{M}\right\}_{c=0}^{M-1}$ contain the same elements of $\left\{O^{(\tilde{c})}_{M}\right\}_{\tilde{c}=0}^{M-1}$, but these are ranked by magnitude, such that $O^{(\tilde{0})}_{M}$ is the greatest output value between the $M$, $O^{(\tilde{1})}_{M}$ the second one, and so on.
  • $\mathbb{P}\left(A\right)$: Denotes the probability associated with event $A$. $\mathbb{P}\left(A\mid B\right)$ indicates the probability of event $A$ given event $B$.
  • $p_{X}\left({x^{\prime}}\right)$: Given a random variable $X$, $p_{X}\left({x^{\prime}}\right)$ denotes the probability density function (p.d.f.) evaluated at $X=x^{\prime}$. Formally, $p_{X}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}}$. For variables with multiple sources of randomness, if some of these sources are either fixed or marginalized, we will specify the active sources of randomness as subscripts. For example, given $X(\chi,\mathcal{W})$, a function of two independent sets of random variables $\chi$ and $\mathcal{W}$, $p_{X}^{(\chi)}\left({x^{\prime}}\right)$ equals $F_{X}^{(\chi)}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}}$.
@@
  • $\mu_{c}\equiv\left\langle{O^{(c)}}\right\rangle$: This notation is a shorthand to represent the mean value of the output node $c$, computed with respect to the dataset, for a fixed set of initialized weights $\mathcal{W}$. In mathematical terms, it denotes the conditional expectation of the output node $c$, given a specific set of weights.
  • $\boldsymbol{\rho}^{(l;m)}$: vector of the $l^{\text{th}}$ hidden layer, $\left(\rho^{(l;m)}_{0},\dots,\rho^{(l;m)}_{\left\lceil\frac{N_{l}-1}{m}\right\rceil}\right)$, after passing through the max pooling layer with kernel size $m$; $\rho^{(l;m)}_{i}$ indicate the component corresponding to node $i$.
  • $\Theta\left(x\right)$: Heaviside step function.
  • $\delta\left(x\right)$: Dirac delta function.

### v5

@@ 5 IGB: emergence and amplification
By leveraging Theorem 4.1, we are able to theoretically estimate the p.d.f. of $G_{0}$, building upon the distributions defined in Eq. (5) and Eq. (6). These estimations, described in App. E.1 and App. G.3.1, enable us to systematically evaluate $f_{G_{0}}\left({g_{0}}\right)$ across a spectrum of model designs.
@@
    $f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)=\mathcal{N}\left(\delta_{o};\Delta_{\mu},2\textrm{Var}_{\chi}\left({O_{c}}\right)\right)\,.$ (11)
@@
    Substituting Eq. (11) into Eq. (4) allows us to invert the equation linking $G_{0}$ to $\Delta_{\mu}$, thereby expressing $\Delta_{\mu}$ as a function of $G_{0}$.
@@
    $f_{\Delta_{\mu}}\left({\delta_{m}}\right)=\mathcal{N}\left(\delta_{m};0,2\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\right).$ (12)
@@
    Starting from Eq. (12) and the relationship $\Delta_{\mu}(G_{0})$ from point (B), we obtain $f_{G_{0}}\left({g_{0}}\right)$ through a change of variables, that is, by applying the formula:
    $f_{G_{0}}\left({g_{0}}\right)\,dg_{0}=f_{\Delta_{\mu}}\left({\delta_{m}(g_{0})}\right)\,d\delta_{m}\,.$ (13)
This analysis not only identifies the architectural elements critical to the emergence of IGB but also quantifies their influence. [^1] We note that, to derive $f_{G_{0}}\left({g_{0}}\right)$, we work in the infinite-width limit. This is however not a necessary condition for the appearance of IGB (see e.g. experiments with finite-width models in App. I), but rather an assumption that allows us to derive the theoretical curves.
@@ 5.1 Single hidden layer
Figure 5: The distribution $f_{G_{0}}\left({g_{0}}\right)$ in a single-hidden-layer perceptron, for different choices of activation functions and with/without max pooling.
Let us begin by considering the case with $L=1$; in this scenario, the element of non-linearity in the network (such as the activation function, pooling layer, etc.) can give rise to IGB. In our analysis we consider three different setups to highlight the different behaviors induced by non-linearity elements. Fig. 5 shows $f_{G_{0}}\left({g_{0}}\right)$ (both empirical histogram and theoretical curves) for the following cases:
  • Linear: $f_{G_{0}}\left({g_{0}}\right)$ asymptotically converges to a delta distribution peaked on $g_{0}=1/2$. [^2] By asymptotically, we mean in the $D\rightarrow\infty$} limit. The theoretical curve shown in the plot takes into account the finite dataset size effects [^3] To take finite dataset-size effects into account, we substitute the population data distribution, with the empirical one, defined over the finite set of $D$ points. (since in real simulations $D<\infty$) as discussed in App. D.
  • ReLU: In this case, $f_{G_{0}}\left({g_{0}}\right)$ does not concentrate at $g_{0}=1/2$ and stays asymptotically wide (we detail this in App. D, Fig. D.1), so $g_{0}$ will, w.h.p., be away from 1/2. The mode of the distribution remains, as for the linear case, at $g_{0}=1/2$. Yet, in this case, the fluctuations from the peak are not due to finite size effects and remain finite in the limit of infinite data points.
@@
    $U^{(1;k)}_{l}$ $=U^{(k)}\left({\left\{H\left({Z^{(1)}_{j}}\right)\right\}_{j\in S_{l}^{k}}}\right)\,$, (14)
    where, in our notation, $U^{(k)}\left({\cdot}\right)$ indicates a pooling function with kernel size equals to $k$, $S_{l}^{k}$ indicates the $l^{\mathrm{th}}$ subgroup of $k$ nodes. In particular, for the max pooling
    $U^{(1;k)}_{l}\equiv\max_{j\in S_{l}^{k}}\left\{\max\left(0,H^{(1)}_{j}\right)\right\}$. (15)
@@
Note how, for the output layer, permutation symmetry in nodes corresponds to symmetry between classes. The latter is always preserved at the ensemble level, [^4] Here, ”ensemble” refers to all possible random initializations specified by the chosen scheme (e.g., Kaiming Normal). in fact in all three cases $\mkern 1.5mu\overline{\mkern-1.5muG_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$, where $\mkern 1.5mu\overline{\mkern-1.5muG_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu\equiv\int f_{G_{0}}\left({g_{0}}\right)g_{0}\,dg_{0}$ indicate the average of $G_{0}\left(\mathcal{W}\right)$ over the ensemble of weight initializations. However, while in the linear case, the symmetry between classes is also conserved on the single element of the ensemble, [^5] In this case, we have $\lim_{D\rightarrow\infty}G_{0}\left(\mathcal{W}\right)=\mkern 1.5mu\overline{\mkern-1.5muG_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$. in the other two cases the single realization diverges, w.h.p. from the mean over the ensemble. NSB thus results in a breaking of the symmetry between classes and a consequent self-averaging breakdown for the observable $G_{0}$. This difference between the two estimates is not due to finite network size effects; our predictions (theoretical curves in Fig. 5) are, in fact, formally exact in the limit of infinite size.
@@
Our results can be extended to a generic activation function. In particular, we can clarify (App. F) what is the fundamental attribute of the activated nodes, i.e. passed through the activation function and the pooling layer, $\{U^{(1;k)}_{i}\}_{i}$, that triggers NSB and consequently IGB:
@@
Consider a Gaussian distributed dataset processed through a single hidden layer perceptron, in the limit of infinite width hidden layer, and whose weights are initialized according to the Kaiming normal scheme (with null bias weights). Then, we will have an absence of IGB in the network if, and only if, for the generic $i^{\text{th}}$ node of the hidden layer: $\left\langle{U^{(1;k)}_{i}}\right\rangle=0\;,\;\;\forall i.$ (16) Or conversely, IGB will emerge if, and only if $\left\langle{U^{(1;k)}_{i}}\right\rangle\neq 0\;,\;\;\forall i.$ (17)
@@ 5.2 Conditions for strong IGB
Consider a centered Gaussian distributed dataset that undergoes preprocessing through the standardization $\psi\left(\boldsymbol{\xi^{(a)}}\right)=\boldsymbol{\xi^{(a)}}+\boldsymbol{K},$ where $\boldsymbol{K}$ is an offset with the same dimensionality of the input. The data is then processed through a Multi-Layer Perceptron (MLP) with $L$ hidden layers. This MLP employs ReLU nonlinearities and max pooling layers with a kernel size of $k$, and the weights are initialized using the Kaiming normal scheme (with zero biases). In the limit of infinite width for hidden layers, the following scenarios are observed: I. Shifting the center of the input distribution by a vector with the norm $|\boldsymbol{K}|$, even when $L=1$ and $k=1$, results in: $\lim_{|\boldsymbol{K}|\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (18) II. With an increase in the kernel size of the pooling layer, even for a single hidden layer ($L=1$), and in the absence of standardization offset ($|\boldsymbol{K}|=0$), one has: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (19) III. In the infinite-depth limit, even in the absence of pooling layers ($k=1$) and standardization ($|\boldsymbol{K}|=0$): $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (20)
@@ 5.3.3 Effects of IGB on Pre-Trained Models
Figure 6: The distribution $f_{G_{0}}\left({g_{0}}\right)$ on pre-trained EfficientNetV2 with the binary dataset (C10). The difference lies in the untrained head: a fully connected layer (PT-ENetV2-FC) versus a MLP (PT-ENetV2-MLP). More details in App. I.4. Increasing head depth amplifies IGB, consistently with our analysis.
@@ Appendix A Notation
$Z^{(l+1)}_{i}$ $=\sum_{j}w^{(l+1)}_{ij}U^{(l;k)}_{j}\,$, (21)
$H^{(l)}_{i}$ $=H\left({Z^{(l)}_{i}}\right)\,$, (22)
$U^{(l;k)}_{i}$ $=U^{(k)}\left({\left\{H\left({Z^{(l)}_{j}}\right)\right\}_{j\in S_{l}^{m}}}\right),$ (23)
[…]
@@
  • $F_{X}\left({x}\right)$: Given a random variable $X$, we denote its cumulative distribution function (c.d.f.) as $F_{X}\left({x}\right)$, where $F_{X}\left({x}\right)=\mathbb{P}\left(X<x\right)$.
  • $F_{X}^{(\boldsymbol{\cdot})}\left({x}\right)$: For random variables depending on multiple sources of randomness, we explicitly indicate the active source of randomness when one is fixed. For example, considering $X(\chi,\mathcal{W})$, a function of two independent sets of random variables $\chi$ and $\mathcal{W}$, $F_{X}^{(\chi)}\left({x}\right)$ denotes the c.d.f. with respect to $\chi$, while holding $\mathcal{W}$ fixed. When $\chi$ has been marginalized, leaving only $\mathcal{W}$ as the active source, the subscript is omitted for conciseness, as there is no ambiguity.
  • $f_{X}\left({x^{\prime}}\right)$: Given a random variable $X$, $f_{X}\left({x^{\prime}}\right)$ denotes the probability density function (p.d.f.) evaluated at $X=x^{\prime}$. Formally, $f_{X}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}}$.
  • $f_{X}^{(\boldsymbol{\cdot})}\left({x^{\prime}}\right)$: For random variables with multiple sources of randomness, we specify the active sources of randomness as subscripts when some are fixed. For example, given $X(\chi,\mathcal{W})$, a function of two independent sets of random variables $\chi$ and $\mathcal{W}$, $f_{X}^{(\chi)}\left({x^{\prime}}\right)$ represents the p.d.f. of $X$ with respect to $\chi$, holding $\mathcal{W}$ fixed. When one of the sources of randomness, such as $\chi$, has been marginalized, there remains a single active source of randomness (e.g., $\mathcal{W}$). In such cases, the subscript can be omitted for simplicity and conciseness, as there is no ambiguity.
  • $G_{c}$: fraction of dataset elements classified as belonging to class $c$. The argument $M$ indicates the total number of output nodes for the variable definition, i.e. the number of classes considered. For binary problems we omit this argument ($G_{0}$) as there is only one non-trivial possibility, i.e. $M=2$ .
  • $\boldsymbol{H}^{(l)}$: vector of the $l^{\text{th}}$ hidden layer nodes, $\left(H^{(l)}_{0},\dots,H^{(l)}_{N_{l}-1}\right)$, after passing through the activation function; $H^{(l)}_{i}$ indicate the component corresponding to node ii.
  • $\boldsymbol{Z}^{(l)}$: vector of the $l^{\text{th}}$ hidden layer nodes, $\left(Z^{(l)}_{0},\dots,Z^{(l)}_{N_{l}-1}\right)$, before passing through the activation function; $Z^{(l)}_{i}$ indicate the component corresponding to node ii.
@@
  • $\mathcal{N}\left(x;\mu,\sigma^{2}\right)$: Given a Gaussian r.v. $X$, we indicate with $\mathcal{N}\left(x;\mu,\sigma^{2}\right)$ the p.d.f. computed at $X=x$, i.e. $\mathcal{N}\left(x;\mu,\sigma^{2}\right)\equiv f_{X}\left({x}\right)=\frac{e^{-\frac{1}{2\sigma^{2}}(x-\mu)^{2}}}{\sqrt{2\pi\sigma^{2}}}$
  • $O_{c,M}$: output layer node; $c$ is the node index; the index M, instead, indicate the total number of nodes considered. For binary problems we omit the subscript index to keep the notation lighter, i.e. $O_{c}$ .
  • $O_{\tilde{0},M}$: the set $\left\{O_{\tilde{c},M}\right\}_{c=0}^{M-1}$ contain the same elements of $\left\{O_{\tilde{c},M}\right\}_{\tilde{c}=0}^{M-1}$, but these are ranked by magnitude, such that $O_{\tilde{0},M}$ is the greatest output value between the $M$, $O_{\tilde{1},M}$ the second one, and so on.
  • $\mathbb{P}\left(A\right)$: Denotes the probability associated with event AA. $\mathbb{P}\left(A\mid B\right)$ indicates the probability of event AA given event $B$.
  • $\boldsymbol{U}^{(l;k)}$: vector of the $l^{\text{th}}$ hidden layer, $\left(U^{(l;k)}_{0},\dots,U^{(l;k)}_{\left\lceil\frac{N_{l}-1}{k}\right\rceil}\right)$, after passing through the max pooling layer with kernel size $k$; $U^{(l;k)}_{i}$ indicate the component corresponding to node ii.
@@
  • $G_{\tilde{c}}$: the set $\left\{G_{c}\right\}_{c=0}^{M-1}$ contain the same elements of $\left\{G_{\tilde{c}}\right\}_{c=0}^{M-1}$, but these are ranked by magnitude, such that $G_{\tilde{0}}$ is the greatest output value between the $M$ elements of the set, $G_{\tilde{1}}$ the second one and so on.
  • $\mu_{c}\equiv\left\langle{O_{c}}\right\rangle$: This notation is a shorthand to represent the mean value of the output node $c$, computed with respect to the dataset, for a fixed set of initialized weights $\mathcal{W}$. In mathematical terms, it denotes the conditional expectation of the output node $c$, given a specific set of weights.
  • $\Theta\left(x\right)$: Heaviside step function. $\Theta\left(x\right)=1\penalty 10000\ \forall x\geq 0$, $\Theta\left(x\right)=0\penalty 10000\ \forall x<0$.
  • $\delta\left(x\right)$: Dirac delta function, defined by the relation $\int_{-\infty}^{\infty}f(x)\delta(x-x_{0})dx=f(x_{0})$.